A shipping box has a volume of 2.5 cubic meters. the box measures 1 meter high by 1.25 meters wide. how long is the box?
Answer by checkley79(3341)

V=L*W*H
2.5=L*1.25*1
1.25L=2.5
L=2.5/1.25
L=2 METERS IS THE LENGTH.
2.5=2*1.25*1
2.5=2.5